OpenStudy (anonymous):

Factor the polyomial completely using any method -4w^8-8w^6+4w^4+8w^2 Medal given for answer

OpenStudy (ybarrap):

-4w^2(w^6 + 2w^4 - 2w^2 - 2) let x = w^2 and solve as before ... need to leave for a while

OpenStudy (anonymous):

Thanks Medal given

OpenStudy (ybarrap):

b back in 1 hr

OpenStudy (mathteacher1729):

http://www.wolframalpha.com/ type factor -4w^8-8w^6+4w^4+8w^2

OpenStudy (ybarrap):

-4w^8-8w^6+4w^4+8w^2 -4w^2(w^6 + 2w^4 - w^2 - 2) let x = w^2 Substitute -4x(x^3 + 2x^2 - x - 2) -4x( x^2(x + 2) - (x + 2) ) -4x( (x^2 - 1) (x + 2) ) -4x ( (x + 1)(x - 1)(x + 2) ) Substitute back -4w^2( (w^2 + 1)(w^2 - 1)(w^2 + 2) ) -4w^2( (w^2 + 1)(w + 1)(w - 1)(w^2 + 2) ) Ans. -4w^2( (w^2 + 1)(w + 1)(w - 1)(w^2 + 2) )